OpenStudy (anonymous):

Please help solve –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

\[-2x +6 \ge 5x-15\] \[-2x + 21 \ge 5x\] \[21 \ge 7x\] \[3 \ge x\]

OpenStudy (anonymous):

ty